John Bowie is trying to determine the amount to set aside so that he will have enough money on hand in 4 years to overhaul the engine on his vintage used car. While there is some uncertainty about the cost of engine overhauls in 4 years, by conducting some research online, John has developed the following estimates.

Engine Overhaul
Estimated Cash Outflow

Probability
Assessment

$270 10%
630 30%
750 50%
870 10%



Click here to view factor tables

How much should John Bowie deposit today in an account earning 8%, compounded annually, so that he will have enough money on hand in 4 years to pay for the overhaul?
Answer #1

Engine Overhaul
Estimated Cash Outflow

(x)

Probability
Assessment

(p)

xp
270 10% 27
630 30% 189
750 50% 375
870 10% 87
total 678

In order to receive 678 $ @ end of 4 years @ 8% compounded annually,

V = a(1+r)n where V=value @ end of n years, a = amount invested, r = rate of int, n =no. of years

678 = a(1+0.08)4

So, A = 498.35.

So, John Bowie should deposit $498.35 today